# A Particle Algorithm for Mean-Field Variational Inference

### v2

@@ 2.2 Key properties of MFVI
For any $d\in\mathbb{Z}_{+}$ and $\mu\in\mathcal{P}_{\mathrm{AC}}(\mathbb{R}^{d})$, we define the entropy of $\mu$ as
$H(\mu)=-\int_{\mathbb{R}^{d}}\mu(\bm{x})\log\mu(\bm{x})\mathrm{d}\bm{x}.$
@@
For any $q=q^{1}\otimes\cdots q^{m}\in\mathcal{Q}$ and $i\in[m]$, define
@@
In other words, for a particle $\boldsymbol{x}$ or a distribution $q$, the subscript or superscript $-i$ means “all components except for the $i$-th one”.
@@
The function $\bar{V}_{i}$ is an approximation (using a single component and a distribution $\mu$) to the potential function $V:\mathbb{R}^{m}\rightarrow\mathbb{R}$. It will be crucial for defining the mean-field particle dynamics. In this paper, any derivative of $\bar{V}_{i}$ is taken with respect to its first argument.
@@
$D_{\mathrm{KL}}(q\|p)=-\mathbb{E}_{\bm{x}\sim q}\log\bigg{(}\frac{\mathrm{d}p}{\mathrm{d}q}(\bm{x})\bigg{)}=\mathbb{E}_{\bm{x}\sim q}V(\bm{x})-H(q)+\log(Z).$
@@
If we minimize the Kullback–Leibler divergence $D_{\mathrm{KL}}(q\|p)$ while keeping $q^{-i}$ fixed, then the optimal $q^{i}$ is clearly ${\mathcal{T}}_{i}(q^{-i})$. Hence, the optimality of $q_{*}=\mathop{\mathrm{argmin}}_{q\in\mathcal{Q}}D_{\mathrm{KL}}(q\|p)$ implies the optimality of the $i$-th component $q_{*}^{i}={\mathcal{T}}_{i}(q_{*}^{-i})$, $\forall i\in[m]$. Thus, we get the following result, which characterizes the optimal $q_{*}$ as the fixed point of a transformation.
@@
$q_{*}$ is the fixed point of transform ${\mathcal{T}}:~{}\mathcal{Q}\to\mathcal{Q}$, $q\mapsto{\mathcal{T}}_{1}(q^{-1})\otimes\cdots\otimes{\mathcal{T}}_{m}(q^{-m})$.
@@ 3 A particle algorithm for mean-field variational inference
In this section, we introduce a particle algorithm for mean-field variational inference, which we term PArticle VI (PAVI). In Section 3.1, we formulate the variational inference problem in terms of Fokker–Planck equations and the corresponding diffusion processes. Then, based on this formulation, we introduce the particle algorithm for practical implementation.
@@ 3.1 Fokker–Planck equations and diffusion processes
According to the general results by Jordan et al. (1998), the gradient flow for the Kullback–Leibler divergence $D_{\mathrm{KL}}(\,\cdot\,\|{\mathcal{T}}_{i}(q^{-i}))$ with respect to the 2-Wasserstein metric is the solution of the Fokker–Planck equation
@@
The Fokker–Planck equation is deeply connected to a stochastic differential equation: the solution $\rho(\,\cdot\,,t)$ of (3.1) is the law of the diffusion process
$\begin{cases}&\mathrm{d}Y_{t}=-f^{\prime}(Y_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},\\ &Y_{0}\sim q^{i},\end{cases}$ (3.2)
where $Y_{t}:[0,\infty)\rightarrow\mathbb{R}$ and $\{B_{t}\}_{t\geq 0}$ is a standard Brownian motion that is independent of $Y_{0}$.
Simultaneous updates of all the marginal distributions $\{q^{i}\}_{i=1}^{m}$, where $q^{-i}$ is fixed for each update, lead to a McKean–Vlasov process:
$\begin{cases}&\mathrm{d}Y_{t}^{i}=-\bar{V}_{i}^{\prime}(Y_{t}^{i},\mathrm{Law}(\bm{Y}_{t}^{-i}))\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t}^{i},\qquad i\in[m],\,t\geq 0\\ &\bm{Y}_{0}\sim q^{1}\otimes\cdots\otimes q^{m}\,,\end{cases}.$ (3.3)
where $\bar{V}_{i}$ is defined as in (2.3), $\bm{Y}_{t}=(Y_{t}^{1},\cdots,Y_{t}^{m})$, and $\{(B_{t}^{1},\cdots,B_{t}^{m})\}_{t\geq 0}$ is an $m$-dimensional standard Brownian motion that is independent of $\bm{Y}_{0}$. The trajectory $t\mapsto\mathrm{Law}(\bm{Y}_{t})$ is the Wasserstein gradient flow for optimizing $D_{\mathrm{KL}}(\,\cdot\,\|p)$ over the mean-field family $\mathcal{Q}$ (Lacker, 2023). Note that the process $\{\bm{Y}_{t}\}_{t\geq 0}$ has independent coordinates because the $m$-dimensional Brownian motion has independent coordinates and the other coordinates $\bm{Y}_{t}^{-i}$ are integrated out in the drift term.
@@ 3.2 PArticle VI
We are ready to present PArticle VI (PAVI), a particle algorithm for mean-field variational inference. To set the stage, we first introduce some convenient notations.
@@
Let $S_{N}$ be the set of all permutations of $[N]$. For any array $\bm{A}\in\mathbb{R}^{m\times N}$, denote by $A^{i,j}$ its $(i,j)$-th entry, and let $\bm{A}^{i,:}=(A^{i,1},\cdots,A^{i,N})$ be the $i$-th row, $\bm{A}^{:,j}=(A^{1,j},\cdots,A^{m,j})$ be the $j$-th column, and $\bm{A}^{-i,j}=(A^{1,j},\cdots,A^{i-1,j},A^{i+1,j},\cdots,A^{m,j})$ be the $j$-th column of $\bm{A}$ with the $i$-th entry removed. For each $i\in[m]$, let $q_{\bm{A}}^{i}=\frac{1}{N}\sum_{j=1}^{N}\delta_{A^{i,j}}$. Define $q_{\bm{A}}=q_{\bm{A}}^{1}\otimes\cdots\otimes q_{\bm{A}}^{m}$.
@@
For each $i\in[m]$, the measure $q^{i}_{\bm{A}}$ is the empirical measure of $\bm{A}^{i,:}$, the $i$-th component for the $N$ particles. $q_{\bm{A}}\in\mathcal{P}(\mathbb{R}^{m})$ is the product measure with $\{q^{i}_{\bm{A}}\}_{i=1}^{m}$ as its marginals. Hence, it has $N^{m}$ atoms in total.
Our first version of PAVI, Algorithm 1, can be viewed as a direct discretization of the McKean–Vlasov process (3.3).
@@
If the columns of $\bm{X}_{0}$ are i.i.d. samples from a probability distribution over $\mathbb{R}^{m}$, $N\to\infty$ and $h\to 0$, then Algorithm 1 becomes the Wasserstein gradient flow in Lacker (2023). Meanwhile, we note that Algorithm 1 is intuitive but impractical: the gradient evaluation is prohibitively expensive due to the cost of averaging over $N^{m-1}$ terms. For computational reasons, we propose a second version of PAVI, presented in Algorithm 2 below, which involves a stochastic approximation of the gradient term.
@@
Algorithm 1 is the limiting case of Algorithm 2 (PAVI) as $B\to\infty$. The finite batch size in PAVI ensures its computational efficiency. The output $\bm{X}_{T}$ from Algorithm 2 is associated with its empirical distribution $q_{\bm{X}_{T}}$, which is a product distribution that approximates $q_{*}$, the solution to (2.2). In the next section, we will prove the convergence of $q_{\bm{X}_{T}}$ towards $q_{*}$ in the $\mathcal{W}_{2}$ distance. Our theory will also provide practical guidance on choosing $B$. It is worth pointing out that convergence guarantees for Algorithm 1 can be obtained by taking the limit $B\to\infty$.
@@ 4 Theoretical analysis
In this section, we provide a theoretical analysis of Algorithm 2, the particle algorithm with stochastic approximation introduced in Section 3. Recall from Section 2.2 that the target distribution $p\in\mathcal{P}_{\mathrm{AC}}(\mathbb{R}^{m})$ has density $Z^{-1}e^{-V(\cdot)}$. We work under the following two regularity assumptions.
@@
Assumption 4.2 (Bounded third-order derivatives).
For any $\bm{x}\in\mathbb{R}^{m}$ and $i\in[m]$, we have $\left|\frac{\partial^{3}}{\partial x_{i}^{3}}V(\bm{x})\right|\leq\widetilde{L}<\infty$.
Under Assumption 4.1, we have the following result from Lacker (2023); Lacker et al. (2024), which shows the existence and uniqueness of the MFVI solution $q_{*}$ and establishes the exponential convergence of the law of $\bm{Y}_{t}$ (recall (3.3)) towards $q_{*}$.
Lemma 4.1 (Theorems 1.1 of Lacker et al. (2024) and Theorem 2.3 of Lacker (2023)).
Under Assumption 4.1, the following holds:
  • (a)
    The optimization problem (2.2) has a unique solution $q_{*}$.
  • (b)
    For any $t\geq 0$, we have
    $\mathcal{W}_{2}(\mathrm{Law}(\bm{Y}_{t}),q_{*})\leq e^{-\alpha t}\mathcal{W}_{2}(\mathrm{Law}(\bm{Y}_{0}),q_{*}).$
Lemma 4.1 focuses on the long-time convergence of the continuous-time process $\bm{Y}_{t}$ but does not provide algorithmic guarantees. Our main theoretical result, presented in Theorem 4.1 below, establishes the first end-to-end theoretical guarantee for an implementable particle algorithm (Algorithm 2) for MFVI. In Theorem 4.1, we provide a non-asymptotic upper bound on the $\mathcal{W}_{2}$ distance between $q_{\bm{X}_{n}}$ (the empirical distribution of the $N$ particles at the $n$-th iteration) and the MFVI solution $q_{*}$.
@@
Suppose that Assumptions 4.1 and 4.2 hold. Consider Algorithm 2 with $N\geq 2$ and $0<h<\min\{2/(\alpha+L),B\alpha/(4L^{2})\}$. For any $T\in\mathbb{Z}_{+}$ and $10\leq n\leq T-$ 1, we have
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{*})}$ $\leq(1-\alpha h/4)^{n}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}$
$\quad+\frac{C\sqrt{m}}{\alpha}\bigg{(}h^{-1}\sqrt{\frac{\log N}{\alpha N}}+\sqrt{h\max\{h,(LB)^{-1}\}(L^{4}/\alpha+\widetilde{L}^{2})}\bigg{)},$
where $C>0$ is a universal constant. In particular, when $B\geq 1/(Lh)$, we have
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{*})}$ $\leq(1-\alpha h/4)^{n}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}+\frac{C\sqrt{m}}{\alpha^{3/2}}\bigg{(}h^{-1}\sqrt{\frac{\log N}{N}}+h\sqrt{L^{4}+\widetilde{L}^{2}\alpha}\bigg{)}.$ (4.1)
Recall that as $B\to\infty$, Algorithm 2 reduces to Algorithm 1. Therefore, the error bound (4.1) also holds for Algorithm 1. The proof of Theorem 4.1 will be presented in Section 5.
By taking $h=1/(LN^{1/4})$ and $B=\lceil 1/(Lh)\rceil$ in Theorem 4.1, we obtain the following corollary.

### v3

@@ 2.2 Key properties of MFVI
For any $d\in\mathbb{Z}_{+}$ and $\mu\in\mathcal{P}_{\mathrm{AC}}(\mathbb{R}^{d})$, we define the entropy of $\mu$ as $H(\mu)=-\int_{\mathbb{R}^{d}}\mu(\bm{x})\log\mu(\bm{x})\mathrm{d}\bm{x}$.
@@
For any $q=q^{1}\otimes\cdots\otimes q^{m}\in\mathcal{Q}$ and $i\in[m]$, define
@@
In words, the notation i-i means “all components except for the $i$-th one.”
@@
The function $\bar{V}_{i}$ is an approximation (using a single component and a distribution $\mu$) to the potential $V:\mathbb{R}^{m}\rightarrow\mathbb{R}$. It will be crucial for defining the mean-field particle dynamics. In this paper, any derivative of $\bar{V}_{i}$ is taken with respect to its first argument.
@@
$D_{\mathrm{KL}}(q\|p)=\mathbb{E}_{\bm{x}\sim q}\log\bigg(\frac{\mathrm{d}q}{\mathrm{d}p}(\bm{x})\bigg)=\mathbb{E}_{\bm{x}\sim q}V(\bm{x})-H(q)+\log(Z).$
@@
If we minimize the Kullback–Leibler divergence $D_{\mathrm{KL}}(q\|p)$ while keeping $q^{-i}$ fixed, then the optimal $q^{i}$ is clearly ${\mathcal{T}}_{i}(q^{-i})$. Hence, the optimality of
$q_{*}\in\mathop{\mathrm{argmin}}_{q\in\mathcal{Q}}D_{\mathrm{KL}}(q\|p)$
implies the optimality of the $i$-th component $q_{*}^{i}={\mathcal{T}}_{i}(q_{*}^{-i})$ for every $i\in[m]$. We immediately obtain the following fixed point characterization.
@@
$q_{*}$ is a fixed point of transform ${\mathcal{T}}:~\mathcal{Q}\to\mathcal{Q}$, $q\mapsto{\mathcal{T}}_{1}(q^{-1})\otimes\cdots\otimes{\mathcal{T}}_{m}(q^{-m})$.
@@ 3 A particle algorithm for mean-field variational inference
In this section, we introduce a particle algorithm for MFVI, which we term PArticle VI (PAVI). In Section 3.1, we formulate the variational inference problem in terms of Fokker–Planck equations and the corresponding diffusion processes. Then, based on this formulation, we introduce the particle algorithm for practical implementation.
@@ 3.1 Fokker–Planck equations and diffusion processes
According to the general results by jordan1998variational, the gradient flow for the Kullback–Leibler divergence $D_{\mathrm{KL}}(\,\cdot\,\|{\mathcal{T}}_{i}(q^{-i}))$ with respect to the 2-Wasserstein metric is the solution to the Fokker–Planck equation
@@
The Fokker–Planck equation is deeply connected to a stochastic differential equation: the solution $\rho(\,\cdot\,,t)$ to (3.1) is the law of the diffusion process
$\begin{cases}&\mathrm{d}Y_{t}=-f^{\prime}(Y_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},\qquad t\geq 0\\ &Y_{0}\sim q^{i},\end{cases}$ (3.2)
where $\{B_{t}\}_{t\geq 0}$ is a standard Brownian motion that is independent of $Y_{0}$. By simultaneously updating all the marginal distributions $\{q^{i}\}_{i=1}^{m}$ according to the above diffusion, we get a McKean–Vlasov process:
[…]
where $\bar{V}_{i}$ is defined as in (2.4), $Y_{t}=(Y_{t,1},\cdots,Y_{t,m})$, and $\{(B_{t,1},\cdots,B_{t,m})\}_{t\geq 0}$ is an $m$-dimensional standard Brownian motion that is independent of $Y_{0}$. The trajectory $t\mapsto\mathrm{Law}(Y_{t})$ is the Wasserstein gradient flow for optimizing $D_{\mathrm{KL}}(\,\cdot\,\|p)$ over the mean-field family $\mathcal{Q}$ (Lac23). Note that the process $\{Y_{t}\}_{t\geq 0}$ has independent coordinates because the $m$-dimensional Brownian motion has independent coordinates and the other coordinates $Y_{t,-i}$ are integrated out in the drift term. It is also worth pointing out that if $Y_{0}\sim q_{*}$, then $Y_{t}\sim q_{*}$ holds for all $t\geq 0$.
@@ 3.2 PArticle VI
We are ready to present the PArticle VI (PAVI), a particle algorithm for mean-field variational inference. To set the stage, we first introduce some convenient notations.
@@
Let $S_{N}$ be the set of all permutations of $[N]$. For any array $\bm{A}\in\mathbb{R}^{m\times N}$, denote by $A^{i,j}$ its $(i,j)$-th entry; let $\bm{A}^{i,:}=(A^{i,1},\cdots,A^{i,N})$ be the $i$-th row, $\bm{A}^{:,j}=(A^{1,j},\cdots,A^{m,j})$ be the $j$-th column, and $\bm{A}^{-i,j}=(A^{1,j},\cdots,A^{i-1,j},A^{i+1,j},\cdots,A^{m,j})$ be the $j$-th column of $\bm{A}$ with the $i$-th entry removed. For each $i\in[m]$, let $q_{\bm{A}}^{i}=\frac{1}{N}\sum_{j=1}^{N}\delta_{A^{i,j}}$. Define $q_{\bm{A}}=q_{\bm{A}}^{1}\otimes\cdots\otimes q_{\bm{A}}^{m}$.
@@
For each $i\in[m]$, $q^{i}_{\bm{A}}$ is the empirical measure of $\bm{A}^{i,:}$. $q_{\bm{A}}\in\mathcal{P}(\mathbb{R}^{m})$ is the product measure with $\{q^{i}_{\bm{A}}\}_{i=1}^{m}$ as its marginals. Hence, it has up to $N^{m}$ atoms.
Our first version of PAVI, Algorithm˜1, can be viewed as a direct discretization of the McKean–Vlasov process (3.3).
@@
If the columns of $\bm{X}_{0}$ are i.i.d. samples from a probability distribution over $\mathbb{R}^{m}$, $N\to\infty$ and $h\to 0$, then Algorithm˜1 becomes the Wasserstein gradient flow in Lac23. Meanwhile, we note that Algorithm˜1 is intuitive but impractical: the gradient evaluation is prohibitively expensive due to the cost of averaging over $N^{m-1}$ terms. For computational reasons, we propose a second version of PAVI, presented in Algorithm˜2 below, which involves a stochastic approximation of the gradient term.
@@
     Draw a random vector $\bm{\xi}_{n}^{i,:}$ from $N(\mathbf{0},\bm{I}_{N})$, the $N$-dimensional standard normal distribution.
@@
Algorithm˜1 is the limiting case of Algorithm˜2 (PAVI) as $B\to\infty$. The finite batch size in PAVI ensures its computational efficiency. The output $\bm{X}_{T}$ from Algorithm˜2 is associated with $q_{\bm{X}_{T}}$, which is a product distribution that approximates the solution $q_{*}$ to (2.2). In the next section, we will prove the convergence of $q_{\bm{X}_{T}}$ towards $q_{*}$ in the $\mathcal{W}_{2}$ distance.
Remark 3 (Block PAVI).
To extend Algorithm˜2 for the block MFVI introduced in Remark˜1, one would use $\bm{X}_{n}\in\mathbb{R}^{(d_{1}+\cdots+d_{m})\times N}$ and operate on the $m$ blocks of coordinates: $X_{n}^{i,j}$ consists of the entries of $\bm{X}_{n}$ in the $(1+\sum_{k=1}^{i-1}d_{k})$-th through $\sum_{k=1}^{i}d_{k}$-th rows and the $j$-th column. The partial derivative $\partial_{i}$ in the function $g_{n}^{i}$ should be replaced with the gradient with respect to the $i$-th block of arguments.
@@ 4 Theoretical analysis
We now provide theoretical guarantees for Algorithm 2, the particle algorithm with stochastic gradient approximation, and its extension to the block case. Recall from Section 2.2 that the target distribution $p\in\mathcal{P}_{\mathrm{AC}}(\mathbb{R}^{m})$ has density $Z^{-1}e^{-V(\cdot)}$. We mainly work under the commonly used assumption below.
@@
The following result shows the existence and uniqueness of the MFVI solution $q_{*}$, and establishes the exponential convergence of the law of $Y_{t}$ in (3.3) towards $q_{*}$.
Lemma 4.1 (Theorems 1.1 of lacker2024mean and Theorem 2.3 of Lac23).
Under Assumption 4.1, the optimization problem (2.2) has a unique solution $q_{*}$, and
$\mathcal{W}_{2}(\mathrm{Law}(Y_{t}),q_{*})\leq e^{-\alpha t}\mathcal{W}_{2}(\mathrm{Law}(Y_{0}),q_{*}),\qquad\forall t\geq 0.$
Lemma 4.1 focuses on the long-time convergence of the continuous-time process $Y_{t}$ but does not provide algorithmic guarantees. Our main theoretical result, presented in Theorem 4.1 below, establishes the first end-to-end theoretical guarantee for an implementable particle algorithm (Algorithm 2) for MFVI. We provide a non-asymptotic upper bound on the $\mathcal{W}_{2}$ distance between $q_{\bm{X}_{n}}$ (the product distribution defined by $N$ particles at the $n$-th iteration) and the MFVI solution $q_{*}$. The proof will be presented in Section 5.
@@
Let Assumption 4.1 hold. Run Algorithm 2 with $N\geq 2$, $0<h<\alpha/(4L^{2})$ and $B\geq 1$. There exists a universal constant $C$ such that for any $n\in\mathbb{N}$,
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{*})}$ $\leq(1-\alpha h/2)^{n/2}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}+C\sqrt{m}\bigg(\sqrt{\frac{\log N}{\alpha N}}+\frac{L\sqrt{h}}{\alpha}\bigg).$
Our non-asymptotic bound quantifies the convergence of PAVI. To the best of our knowledge, this is the first such guarantee for particle-based MFVI. The error comes from four sources:
  • Initialization: The impact of the initial discrepancy is bounded by $(1-\alpha h/2)^{n/2}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}$, which decays exponentially over time.
  • Particle approximation: The error of discretizing the optimal solution $q_{*}$ by $N$ particles is $O(\sqrt{\frac{m\log N}{\alpha N}})$. This is sharp up to a logarithmic factor: if $N\geq 3$ and $\bm{Y}\in\mathbb{R}^{m\times N}$ has independent columns drawn from $q=N(\mathbf{0},\alpha^{-1}\bm{I}_{m})$, then $\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q)}\geq\sqrt{\frac{cm\log\log N}{\alpha N}}$ holds with some universal constant $c$. See Section˜5.2 for a proof.
  • Time discretization: The finite difference (Euler-Maruyama) scheme incurs an error that is $O(\alpha^{-1}L\sqrt{mh})$.
  • Stochastic gradient approximation: Algorithm 2 uses $B$ random samples to estimate the gradient in Algorithm 1. For any $B\geq 1$, our proof shows that the stochastic error is $O(\alpha^{-1}L\sqrt{mh})$. This is dominated by the time discretization error and hence does not appear explicitly in the final inequality. One can safely use $B=1$ for computational efficiency.
Our theory can be generalized to the block PAVI algorithm discussed in Remark˜3. Suppose that the target distribution $p\in\mathcal{P}_{\mathrm{AC}}(\mathbb{R}^{d})$ has density $Z^{-1}e^{-V(\cdot)}$, and $\alpha\bm{I}_{d}\preceq\nabla^{2}V\preceq L\bm{I}_{d}$ holds for some $0<\alpha\leq L<\infty$. Let $\{d_{i}\}_{i=1}^{m}$ be positive integers that sum up to $d$. The variational problem (2.2) over the block mean-field family (2.3) still has a unique solution $q_{*}=q_{*}^{1}\otimes\cdots\otimes q_{*}^{m}$ (Lac23). We have the following error bound. The proof can be found in Section˜5.3.
Theorem 4.2.
Under the above assumptions, the block version of Algorithm 2 with $N\geq 2$, $0<h<\alpha/(4L^{2})$ and $B\geq 1$ satisfies
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{*})}$ $\leq(1-\alpha h/2)^{n/2}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}+\sqrt{\sum_{i=1}^{m}\mathbb{E}\mathcal{W}_{2}^{2}(\hat{q}_{N}^{i},q_{*}^{i})}+\frac{4L\sqrt{hd}}{\alpha},$
where $\hat{q}_{N}^{i}$ is the empirical distribution formed by $N$ independent samples from $q_{*}^{i}$.
When $d_{1}=\cdots=d_{m}=1$, standard concentration inequality of univariate empirical measure (e.g., Lemma˜6.5) shows that $\mathbb{E}\mathcal{W}_{2}^{2}(\hat{q}_{N}^{i},q_{*}^{i})=O(\frac{\log N}{\alpha N})$. Then, Theorem˜4.2 reduces to Theorem˜4.1. For general $d_{i}$’s, one can bound $\mathbb{E}\mathcal{W}_{2}^{2}(\hat{q}_{N}^{i},q_{*}^{i})$ using the results in FGu15.
We now come back to the original version of Algorithm˜2. Theorem˜4.1 suggests taking $B=1$ for computational efficiency, and making $\frac{L}{\alpha}\sqrt{h}$ and $\sqrt{\frac{\log N}{\alpha N}}$ have the same order. Based on these observations, we get the following user-friendly error bound. See Section˜5.4 for the proof.